# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using AMPT

### v1

@@ Abstract
The study of symmetry plane correlations (SPCs) can be useful in characterizing the direction of the anisotropic emission of produced particles in the final state. The study of SPCs provides an independent method to understand the transport properties of the system formed in heavy-ion collisions. Similar to anisotropic flow coefficients, which are largely influenced by the initial spatial anisotropy, SPCs also depend upon the participant plane correlation measured using the participating nucleons of the collision overlap region. In this paper, SPCs have been studied in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using the event generator AMPT. In addition to their behaviour with the changing centrality of the collision, their event shape dependence has also been studied for the first time, using the event shape classifier transverse spherocity. The Gaussian estimator has been used to evaluate the correlations, and these have been compared to the participant plane correlations defined in an analogous way to the symmetry plane correlations and a qualitative match has been found between them. These event-shape differentiated symmetry plane correlations can be used to deduce the presence of higher-order anisotropies in the initial energy distribution, thus giving insight into the initial geometry of the colliding system, among other applications like model development and model testing using Bayesian analyses.
@@ I Introduction
In the heavy-ion collisions at ultra-relativistic speeds, the formation of a deconfined state of partons (quarks and gluons), called the quark-gluon plasma (QGP), is anticipated [1, 2, 3]. QGP is transient in nature, having a lifetime of the order of $10^{-23}$ seconds [4]. Thus, presently, QGP formation can only be studied through the study of various indirect signatures such as collectivity [6, 5], jet quenching [7], charmonia suppression [8], strangeness enhancement [9] to name a few. QGP is theorized to behave like a perfect fluid. Studies based on anisotropic flow at the RHIC and LHC energies suggest that the ratio of specific shear viscosity to entropy density ($\eta_{s}/s$) is close to $1/4\pi$ which is the lower bound of $\eta_{s}/s$ imposed by AdS/CFT calculations [10, 11, 12, 13, 14, 15, 16]. Further, for a smaller value of $\eta_{s}/s$, the hydrodynamic nature of the QGP can transform the initial state spatial anisotropy of the collision overlap region defined by the participant nucleons into the final state momentum space azimuthal anisotropy of the produced charged particles [17, 5, 18]. The final state azimuthal anisotropy of the charged particles can be quantified using the coefficients of the Fourier expansion of azimuthal distributions of the particles, as follows [18].
@@
Here, $\phi$ denotes the azimuthal angle in the momentum space. The Fourier coefficients, $v_{n}$, are referred to as the anisotropic flow coefficients, and the parameters $\psi_{n}$ are called the symmetry plane angles. Together, these observables give us an idea of how the particles have been ejected. For example, the second order flow coefficient $v_{2}$ is called the elliptic flow [19] and quantifies the extent up to which the charged particle distribution in the ($p_{x}-p_{y}$) axis can be described by an ellipse. The anisotropic flow coefficients can serve as an instrument to probe for the initial state geometry of the colliding nuclei due to the collective hydrodynamic expansion of the QGP medium [20]. Further discussions regarding the mathematical properties of the flow decomposition of the transverse momentum distribution of the final state particles can be found in Ref. [21].
Some trivial properties of the anisotropic flow coefficients and the symmetry plane angles corresponding to a particular harmonic can be directly inferred from Eq. (1) using the underlying fact that if a change of azimuthal coordinates is made from $\phi\to-\phi$, then the distribution $f(\phi)$, in Eq. (1) remains invariant. Due to this fact, if the Fourier sum in Eq. (1) was to be re-expressed in terms of complex exponential terms, then $v_{n}=v_{-n}$ and $\psi_{n}=\psi_{-n}$. A study of the symmetry plane angles by itself is not feasible as a function of centrality or other event classifiers. This is because, in each collision, there is a random orientation of the reaction plane [22], hence changing the symmetry planes randomly as well. This leads to an event average, $\langle\psi_{n}\rangle=0$. Thus, instead of explicitly studying $\psi_{n}$, it is reasonable to study the correlation between the symmetry plane angles with different event classifiers, including collision centrality. Similar to anisotropic flow coefficients, the correlations of the symmetry plane angles of different orders are sensitive to both initial state correlations and the transport properties of the system. Thus, a study of symmetry plane angles can provide an independent method to understand the initial state transport properties of the medium formed in heavy-ion collisions [23]. In this direction, using the definition of the Gaussian estimator for the symmetry plane correlations (SPC), the correlation among the symmetry planes of different orders has been studied. The computation of symmetry plane angles and of anisotropic flow coefficients involving multi-particle correlations is given below[21]:
@@
In the above equation, the LHS denotes the combination of anisotropic flow coefficients ($v_{n}$) and symmetry plane angles ($\psi_{n}$). The values of the harmonics $n_{i}$ and indices $a_{i}$ are such that $a_{i}$ are positive and hence denote the number of times that a flow term of order $n_{i}$, or equivalently $-n_{i}$ occurs. There is kept independence in the choice of harmonics, which is necessary to define different kinds of correlations appropriately [24], such that the effect of random orientation of the reaction plane from event to event, and the effect of trivial periodicity, i.e., Eq. (1) being invariant under the transformation $\psi_{n}\to\psi_{n}+\frac{2m\pi}{n}$, for integers $0\leq m<n$ play no role when an event average of the symmetry plane correlations is computed. The coefficients are to be chosen so that:
@@
The above choice of coefficients ensures that the two caveats necessary to construct a meaningful physical observable to estimate the correlation between different symmetry planes are taken care of. The integer $l$ in the RHS of Eq. (I) denotes the number of particles that need to be correlated in order to compute the LHS. The particle’s final state azimuthal angles $(\phi_{1},\phi_{2},...,\phi_{l})$ are such that all possible multiplets of particles, with appropriate cuts (Section II.3) are considered and their associated harmonics $(n_{1},n_{2},...,n_{l})$ are the same as harmonics $n_{i}$ which appear in the LHS of Eq. (I), $a_{i}$ times. That is, the correlation in the RHS of Eq. (I) is computed using $a_{i}$ different particles associated with the harmonics $n_{i}$.
@@
The collective behaviour, which occurs due to hydrodynamic expansion of the QGP medium, is expected to be sensitive to the initial conditions of the system of the two colliding nuclei [25]. Hence, the various flow coefficients and symmetry planes could give access to the initial configuration of the nuclei if the system has a linear response to its initial conditions. More explicitly, in such a scenario, the SPCs should exhibit similar behaviour to the initial state, participant plane correlations (PPCs), with changing centrality. Further, due to event-by-event fluctuation in the geometry of the collision overlap region, both initial eccentricity and participant plane angle can vary event-by-event. Thus, in a particular centrality class, both anisotropic flow coefficients and symmetry plane angles can have a broad range distribution. This can lead to changes in symmetry plane correlations within a centrality bin. These varying magnitudes of symmetry plane correlations can be studied using event-shape observables. Transverse spherocity [26, 27, 28] is one such event shape observable which can identify the isotropic events, dominated with soft interactions from the jetty events dominated with hard interactions. Studies based on a multi-phase transport model show the applicability of transverse spherocity to separate events in heavy-ion collisions at the LHC energies [32, 31, 30, 29, 20]. The necessity of studying SPCs by classifying collision events by their shapes is obvious since it could give us insight into the variation of the SPCs with the processes that lead to final particle production with a certain geometry. Also, some correlations, for example, which have been experimentally found to be zero, e.g., the correlation of $\psi_{2}$ and $\psi_{3}$ [23], can be verified if they are always zero or happen to be so due to the averaging over events of all shapes, and actually have non-zero correlation in specific event shapes. Such a study will give insight into the usability of different event shapes to probe into correlations amongst symmetry planes. The event shape classifier used for this purpose is transverse spherocity (hereafter referred to as spherocity or $S_{0}$ for brevity).
@@
The paper is organized as follows. We start with a brief introduction in Section I, followed by the description of the methodology of event generation, i.e., the details of the AMPT parameters and the phase space cuts on particle selection for SPC evaluation in Section II. The results are discussed in Section III. Finally, the paper is summarised in Section IV.
@@ II.1 A Multi-phase Transport model (AMPT)
A multi-phase transport model (AMPT) [33] is a Monte Carlo (MC) based transport model that simulates heavy-ion collisions by explicitly dealing with non-equilibrium dynamics of the collision system. AMPT involves four major stages. In the string melting version of AMPT (AMPT-SM), the collision between two nuclei is initialized using HIJING [34, 35], followed by partonic scatterings calculated using Zhang’s Parton Cascade (ZPC) [36], followed by hadronization using quark coalescence model [37], and finally evaluating the hadronic rescatterings using a relativistic transport (ART) framework [38, 39]. In this study, we use AMPT to simulate Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5020$ GeV. The parameters in the Lund symmetric splitting function are set to $a=0.30$ and $b=0.15$ [40]. The parton screening mass is set to 1.2408 $\rm fm^{-1}$ [36]. Flags for shadowing, initial and final state radiations, and random orientation of the reaction plane angle were kept ‘on’. A total of 110,000 minimum bias Pb-Pb collision events are generated for this study using the settings mentioned above.
@@ II.2 Transverse spherocity
Transverse spherocity, which is used for event shape classification, is defined as follows [20, 41, 42]:
$S_{0}=\frac{\pi^{2}}{4}\min_{\hat{n}}\left[\left(\frac{\Sigma_{i}|\vec{p}_{T_{i}}\times\hat{n}|}{\Sigma_{i}|\vec{p}_{T_{i}}|}\right)^{2}\right]$ (4)
@@
Where $\vec{p}_{T_{i}}$ refers to the transverse momentum vector of the charged hadron track $i$ in an event and $\hat{n}$ is a parameter unit vector in the plane transverse to the beam axis which is used to minimize the quantity within the square brackets, and the constant factor at the front is for the purpose of normalization. In accordance with its definition in Ref. [20], only those charged hadronic tracks with transverse momenta $p_{\rm T}>0.15$ GeV/c have been used for the evaluation of spherocity. An additional constraint that has been used is that only those events have been considered for this study, in which there are at least five charged hadrons satisfying the given transverse momentum cut. The extreme limits of $S_{0}\to 0$ and $S_{0}\to 1$ correspond to the event shape being pencil-like and isotropic, respectively. This is mainly because, in a pencil-like event, the particles emerge collimated along a particular direction, and hence, aligning $\hat{n}$ in that direction, all the terms in the numerator of Eq. (4) vanish, leading to $S_{0}\rightarrow 0$. Using the reverse arguments, we have that for an isotropic event, $S_{0}\rightarrow 1$. However, for all events, their $S_{0}$ values lie in between these extreme limits. Further details regarding spherocity can be found in Ref. [20].
II.3 Event and track selection

### v2

# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using a multiphase transport model

@@ Abstract
The study of symmetry plane correlations (SPCs) can be useful in characterizing the direction of the anisotropic emission of produced particles in the final state. The study of SPCs provides an independent method to understand the transport properties of the system formed in heavy-ion collisions. Similar to anisotropic flow coefficients, which are largely influenced by the initial spatial anisotropy, SPCs also depend upon the participant plane correlations measured using the participating nucleons of the collision overlap region. In this paper, SPCs have been studied in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using the event generator AMPT. In addition to their behaviour with the changing centrality of the collision, their event shape dependence has also been studied for the first time, using the event shape classifier transverse spherocity. The Gaussian estimator has been used to evaluate the correlations, and these have been compared to the participant plane correlations defined in an analogous way to the symmetry plane correlations, and a qualitative match has been found between them. These event-shape differentiated symmetry plane correlations can be used to deduce the presence of higher-order anisotropies in the initial energy distribution, thus giving insight into the initial geometry of the colliding system, among other applications like model development and model testing using Bayesian analyses.
@@ I Introduction
In the heavy-ion collisions at ultra-relativistic speeds, the formation of a deconfined state of partons (quarks and gluons), called the quark-gluon plasma (QGP), is anticipated [1, 2, 3]. QGP is transient in nature, having a lifetime of the order of $10^{-23}$ seconds [4]. Thus, presently, QGP formation can only be studied through the study of various indirect signatures such as collectivity [6, 5], jet quenching [7], charmonia suppression [8], strangeness enhancement [9], to name a few. QGP is theorized to behave like a perfect fluid. Studies based on anisotropic flow at the RHIC and LHC energies suggest that the ratio of specific shear viscosity to entropy density ($\eta_{s}/s$) is close to $1/4\pi$ which is the lower bound of $\eta_{s}/s$ imposed by AdS/CFT calculations [10, 11, 12, 13, 14, 15, 16]. Further, for a smaller value of $\eta_{s}/s$, the hydrodynamic nature of the QGP can transform the initial state spatial anisotropy of the collision overlap region defined by the participant nucleons into the final state momentum space azimuthal anisotropy of the produced charged particles [17, 5, 18]. The final state azimuthal anisotropy of the charged particles can be quantified using the coefficients of the Fourier expansion of azimuthal distributions of the particles, as follows [18].
@@
Here, $\phi$ denotes the azimuthal angle in the momentum space. The Fourier coefficients, $v_{n}$, are referred to as the anisotropic flow coefficients, and the parameters $\psi_{n}$ are called the symmetry plane angles. Together, these observables give us an idea of how the particles have been ejected. For example, the second order flow coefficient $v_{2}$ is called the elliptic flow [19] and quantifies the extent up to which the charged particle distribution in the ($p_{x}-p_{y}$) axis can be described by an ellipse. The anisotropic flow coefficients can serve as an instrument to probe the initial state geometry of the colliding nuclei due to the collective hydrodynamic expansion of the QGP medium [20]. Further discussions regarding the mathematical properties of the flow decomposition of the transverse momentum distribution of the final state particles can be found in Ref. [21].
Some trivial properties of the anisotropic flow coefficients and the symmetry plane angles corresponding to a particular harmonic can be directly inferred from Eq. (1) using the underlying fact that if a change of azimuthal coordinates is made from $\phi\to-\phi$, then the distribution $f(\phi)$, in Eq. (1) remains invariant. Due to this fact, if the Fourier sum in Eq. (1) were to be re-expressed in terms of complex exponential terms, then $v_{n}=v_{-n}$ and $\psi_{n}=\psi_{-n}$. A study of the symmetry plane angles by itself is not feasible as a function of centrality or other event classifiers. This is because, in each collision, there is a random orientation of the reaction plane [22], hence changing the symmetry planes randomly as well. This leads to an event average, $\langle\psi_{n}\rangle=0$. Thus, instead of explicitly studying $\psi_{n}$, it is reasonable to study the correlation between the symmetry plane angles with different event classifiers, including collision centrality. Similar to anisotropic flow coefficients, the correlations of the symmetry plane angles of different orders are sensitive to both initial state correlations and the transport properties of the system. Thus, a study of symmetry plane angles can provide an independent method to understand the initial state transport properties of the medium formed in heavy-ion collisions [23]. In this direction, using the definition of the Gaussian estimator for the symmetry plane correlations (SPC), the correlation among the symmetry planes of different orders has been studied. The computation of symmetry plane angles and of anisotropic flow coefficients involving multi-particle correlations is given below[21]:
@@
In the above equation, the LHS denotes the combination of anisotropic flow coefficients ($v_{n}$) and symmetry plane angles ($\psi_{n}$). The values of the harmonics $n_{i}$ and indices $a_{i}$ are such that $a_{i}$ are positive and hence denote the number of times that a flow term of order $n_{i}$, or equivalently $n_{i}$ occurs. There is kept independence in the choice of harmonics, which is necessary to define different kinds of correlations appropriately [24], such that the effect of random orientation of the reaction plane from event to event, and the effect of trivial periodicity, i.e., Eq. (1) being invariant under the transformation $\psi_{n}\to\psi_{n}+\frac{2m\pi}{n}$, for integers $0\leq m<n$ play no role when an event average of the symmetry plane correlations is computed. The coefficients are to be chosen so that:
@@
The above choice of coefficients ensures that the two caveats necessary to construct a meaningful physical observable to estimate the correlation between different symmetry planes are taken care of. The integer $l$ in the RHS of Eq. (I) denotes the number of particles that need to be correlated in order to compute the LHS. The particle’s final state azimuthal angles $(\phi_{1},\phi_{2},...,\phi_{l})$ are such that all possible multiplets of particles, with appropriate cuts (Section II.4) are considered and their associated harmonics $(n_{1},n_{2},...,n_{l})$ are the same as harmonics $n_{i}$ which appear in the LHS of Eq. (I), $a_{i}$ times. That is, the correlation in the RHS of Eq. (I) is computed using $a_{i}$ different particles associated with the harmonics $n_{i}$.
@@
The collective behaviour, which occurs due to hydrodynamic expansion of the QGP medium, is expected to be sensitive to the initial conditions of the system of the two colliding nuclei [25]. Hence, the various flow coefficients and symmetry planes could give access to the initial configuration of the nuclei if the system has a linear response to its initial conditions. More explicitly, in such a scenario, the SPCs should exhibit similar behaviour to the initial state, participant plane correlations (PPCs), with changing centrality. Further, due to event-by-event fluctuation in the geometry of the collision overlap region, both initial eccentricity and participant plane angle can vary event-by-event. Thus, in a particular centrality class, both anisotropic flow coefficients and symmetry plane angles can have a broad range distribution. This can lead to changes in symmetry plane correlations within a centrality bin. These varying magnitudes of symmetry plane correlations can be studied using event-shape observables. Transverse spherocity [26, 27, 28] is one such event shape observable that can identify the isotropic events, dominated by soft interactions, from the jetty events dominated by hard interactions. Studies based on a multi-phase transport model show the applicability of transverse spherocity to separate events in heavy-ion collisions at the LHC energies [32, 31, 30, 29, 20, 33, 34].
The necessity of studying SPCs by classifying collision events by their shapes is obvious since it could give us insight into the variation of the SPCs with the processes that lead to final particle production with a certain geometry. Also, some correlations, for example, which have been experimentally found to be zero, e.g., the correlation of $\psi_{2}$ and $\psi_{3}$ [23], can be verified if they are always zero or happen to be so due to the averaging over events of all shapes, and actually have non-zero correlation in specific event shapes. Such a study will provide insight into the usability of different event shapes to probe into correlations amongst symmetry planes. The event shape classifier used for this purpose is transverse spherocity (hereafter referred to as spherocity or $S_{0}$ for brevity). Previously, the event shape dependence of SPCs involving only two planes has been studied, for example, in Ref. [35, 36]. However, these studies did not use the Gaussian Estimator to evaluate the SPCs; hence, they were biased by the magnitude of the involved anisotropic flow coefficients themselves, as was pointed out in Ref. [37].
@@
The paper is organized as follows. We start with a brief introduction in Section I, followed by the description of the methodology of event generation, i.e., the details of a multi-phase transport model (AMPT) parameters and the phase space cuts on particle selection for SPC evaluation in Section II. The results are discussed in Section III.  Finally, the paper is summarised in Section IV.
@@ II.1 A Multi-phase Transport model (AMPT)
AMPT [38] is a Monte Carlo (MC) based transport model that simulates heavy-ion collisions by explicitly dealing with non-equilibrium dynamics of the collision system. AMPT involves four major stages. In the string melting version of AMPT (AMPT-SM), the collision between two nuclei is initialized using HIJING [39, 40], followed by partonic scatterings calculated using Zhang’s Parton Cascade (ZPC) [41], followed by hadronization using quark coalescence model [42], and finally evaluating the hadronic rescatterings using a relativistic transport (ART) framework [43, 44]. In this study, we use AMPT to simulate Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5020$ GeV. The parameters in the Lund symmetric splitting function are set to $a=0.30$ and $b=0.15$ [45]. The parton screening mass is set to 1.2408 $\rm fm^{-1}$ [41]. Flags for shadowing, initial and final state radiations, and random orientation of the reaction plane angle were kept ‘on’. A total of 210,000 minimum bias Pb-Pb collision events are generated for this study using the settings mentioned above.
The $p_{\rm T}$-integrated anisotropic flow coefficients, such as $v_{2}$, $v_{3}$, $v_{4}$, in the string melting version of AMPT, which is used in the present study, are shown to slightly overestimate the experimental measurements [29, 20, 33]. Additionally, AMPT-SM overestimates the $p_{\rm T}$-differential $v_{2}$ in low to intermediate $p_{\rm T}$ regions (coalescence) and underestimates in the high-$p_{\rm T}$ region [46, 47]. AMPT has been shown to qualitatively reproduce many of the experimental features of anisotropic flow, as shown in the previous references, which makes it a good choice to study the symmetry plane correlations.
@@ II.2 Transverse spherocity
Transverse spherocity, which is used for event shape classification, is defined as follows [20, 48, 49, 34]:
$S_{0}=\frac{\pi^{2}}{4}\min_{\hat{n}}\left[\left(\frac{\Sigma_{i}|\vec{p}_{{\rm T}_{i}}\times\hat{n}|}{\Sigma_{i}|\vec{p}_{{\rm T}_{i}}|}\right)^{2}\right]$ (4)
@@
Where $\vec{p}_{{\rm T}_{i}}$ refers to the transverse momentum vector of the charged hadron track $i$ in an event and $\hat{n}$ is a unit vector in the plane transverse to the beam axis which is used to minimize the quantity within the square brackets, and the constant factor $\pi^{2}/4$, at the front is for normalization to unity. In accordance with its definition in Ref. [20], only those charged hadronic tracks with transverse momenta $p_{\rm T}>0.15$ GeV/c have been used for the evaluation of spherocity. An additional constraint that has been used is that only those events have been considered for this study in which there are at least five charged hadrons satisfying the given transverse momentum cut. The extreme limits of $S_{0}\to 0$ and $S_{0}\to 1$ correspond to the event shape being pencil-like and isotropic, respectively. This is mainly because, in a pencil-like event, the particles emerge collimated along a particular direction, and hence, aligning $\hat{n}$ in that direction, all the terms in the numerator of Eq. (4) vanish, leading to $S_{0}\rightarrow 0$. Using the reverse arguments, we have that for an isotropic event, $S_{0}\rightarrow 1$. However, for all events, their $S_{0}$ values lie in between these extreme limits. Further details regarding spherocity can be found in Ref. [20].
II.3 Participant and symmetry plane correlations
From Eq. (I), we can reckon, for a single event, $v_{n}e^{in\psi_{n}}=\langle e^{in\phi}\rangle$. Now, as Pb-Pb collisions at the LHC are associated with high multiplicity, therefore, from event to event, $Re(v_{n})$ and $Im(v_{n})$ (where Re(z),Im(z)Re(z),~Im(z) denote the real and imaginary part of a complex number $z$ respectively) are both expected to have a Gaussian distribution in the Argand plane [37]. Using this assumption, the correlation between symmetry planes, which are independent of the effect of flow coefficients, has been derived in Ref. [37]. The Gaussian estimator for the symmetry plane correlations is given by [23]:
$\langle\cos(a_{1}n_{1}\psi_{n_{1}}+a_{2}n_{2}\psi_{n_{2}}+...+a_{k}n_{k}\psi_{n_{k}})\rangle_{\rm GE}=\sqrt{\frac{\pi}{4}}\frac{\langle{v_{n_{1}}^{a_{1}}v_{n_{2}}^{a_{2}}...v_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\psi_{n_{1}}+a_{2}n_{2}\psi_{n_{2}}+...+a_{k}n_{k}\psi_{n_{k}})}\rangle}{\sqrt{\left\langle{v_{n_{1}}^{2a_{1}}v_{n_{2}}^{2a_{2}}...v_{n_{k}}^{2a_{k}}}\right\rangle}}$ (5)
Here, the angular brackets denote event averaging over a suitable set of events in a particular event class (detailed in Table 1). Further, following its definition in Ref. [37], and as explained in Eq. (3), the choice of harmonics and their powers is made such that $\sum_{i=1}^{k}a_{i}n_{i}=0$. An example of its use and evaluation using multi-particle correlations can be found in Ref. [23].
In the LHS of this equation, $\epsilon_{n}$ are eccentricities of the participating nucleons in the collision overlap region, and $\Phi_{n}$, their corresponding symmetry angles, are called the participant plane angle. The RHS shows the way to compute the combination of $\epsilon_{n}$’s and $\Phi_{n}$’s. The number of particles that are involved in computing the multi-particle correlations in the RHS is $\sum_{i=1}^{k}a_{i}=l$. In the RHS, the angles $\tilde{\phi}_{i}$’s are the azimuthal angles in the coordinate space of the wounded nucleons. Correspondingly, $r_{i}$, which appear to the power $n_{i}$ in the weight factors, are the radial coordinates in the plane transverse to the beam axis. The index $i$ in $r_{i}$ indicates that the radial coordinate being considered is that of the wounded nucleon whose azimuthal angle is $\tilde{\phi}_{i}$. It is important to note that Eq. (II.3) does not entail the correct definition of $\epsilon_{1}e^{i\Phi_{1}}$[50], as well as it fails to account for the minus signs when it comes to computing just one harmonic order of initial state anisotropies. The absence of these features, however, is not an issue for the study of participant plane correlations, as first-order harmonic has not been considered for participant correlations in this paper, and in addition to that, the choice of harmonics $n_{i}$ would lead to all the negative signs canceling when we account for the rotational invariance of the PPCs under the random orientation of the reaction plane from event to event. In case one wishes to study the participant plane correlations involving the first harmonic, i.e., $n=1$, an explicit $r^{3}$ weight factor must be used [50]. Finally, in analogy to Eq. (5), we have the following definition of the Gaussian estimator of the PPCs:
$\langle\cos(a_{1}n_{1}\Phi_{n_{1}}+a_{2}n_{2}\Phi_{n_{2}}+...+a_{k}n_{k}\Phi_{n_{k}})\rangle_{\rm GE}=\sqrt{\frac{\pi}{4}}\frac{\langle{\epsilon_{n_{1}}^{a_{1}}\epsilon_{n_{2}}^{a_{2}}...\epsilon_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\Phi_{n_{1}}+a_{2}n_{2}\Phi_{n_{2}}+...+a_{k}n_{k}\Phi_{n_{k}})}\rangle}{\sqrt{\left\langle{{\epsilon_{n_{1}}^{2a_{1}}\epsilon_{n_{2}}^{2a_{2}}...\epsilon_{n_{k}}^{2a_{k}}}}\right\rangle}}$ (7)
II.4 Event and track selection